# Tomography for Plasma Imaging: a Unifying Framework for Bayesian Inference

### v2

@@ 6 Summary, Conclusion and Next Steps
In this work, we discussed a unifying framework for Bayesian inference from plasma imaging tomographic data. We showed that many popular plasma emission tomographic techniques (such as isotropic/anisotropic approaches based on generalized Tikhonov or MFI regularization, Gaussian Process Tomography, and Maximum Likelihood tomography) can be naturally understood within a common Bayesian framework. Operating within such a framework, we leveraged well-studied optimization and inference concepts to propose a versatile pipeline that relies on simple, robust, and efficient algorithms. We applied this pipeline to the problem of SXR imaging at TCV, demonstrating that the proposed approach enables principled, useful inference from noisy, sparse-view tomographic data. In particular, we validated the pipeline by conducting a phantom-based study over a large set of model phantoms. We showed that MAP computation can be complemented by MCMC methods for posterior sampling, which allow complex statistical analysis but are more computationally expensive; such a complementary vision, with frameworks combining quick MAP-based analyses with detailed MCMC-based computations on data of particular interest, is becoming increasingly popular. Finally, we highlighted some of the inherent limitations of sparse-view tomography and the crucial role played by the prior in such an underdetermined setting: careful prior modeling is clearly essential. Open access is provided to the computational routines implemented in the context of this work, in an effort towards reproducibility and reusability: the link to the project repository can be found at the beginning of Section 5.
@@
For future work, we identify an applied and a methodological axis. From the application viewpoint, we will assess the scientific potential of the proposed pipeline by its application to tomographic reconstruction from the widest possible range of experimental TCV data. In particular, we will broaden the scope from SXR imaging to include bolometry and, potentially, visible-light imaging applications. This can be readily achieved within the generality of the proposed inference framework. From the methodological and modeling points of view, we plan to investigate priors that are different from the classical smoothness approaches traditionally considered for plasma imaging. Specifically, we intend to consider sparsity-promoting priors [32, 46] and recently popularized deep learning-based priors [62].
@@ A Phantom Analysis: supplementary results
$N_{1}$ $0.9\!\cdot\!10^{-2}$ $0.9\!\cdot\!10^{-2}$ $0.6\%$ $1.0\%$ $3.7\,cm$ $4.4\%$ $1.1\%$
$N_{2}$ $1.3\!\cdot\!10^{-2}$ $1.3\!\cdot\!10^{-2}$ $1.1\%$ $1.9\%$ $4.3\,cm$ $5.9\%$ $1.6\%$
$N_{3}$ $1.2\!\cdot\!10^{-2}$ $1.3\!\cdot\!10^{-2}$ $0.9\%$ $1.4\%$ $4.6\,cm$ $5.1\%$ $1.7\%$
@@ B Handling Non-Smooth Terms: Proximal Operators and Moreau-Yosida Approximations
$-\log p(\mathbf{x}|\mathbf{y})=f(\mathbf{x})+g(\mathbf{x})\;,$ (26)
@@
In this work, we employ the reconstruction model in Eq.(24). While the exponential terms in Eq.(24) are smooth, the positivity constraint prior $\mathcal{P}_{\mathbf{x}\geq 0}=\exp(-\mathbbm{1}_{\mathbf{x}\geq 0})$ leads to a non-smooth, proximable term in the negative log-posterior. Therefore, we split $-\log p(\mathbf{x}|\mathbf{y})$ into a smooth and non-smooth component as in Eq.(26), with $g=\mathbbm{1}_{\mathbf{x}\geq 0}$ defined as in Eq.(28) for $\mathcal{C}=\{\mathbf{x}\in\mathbb{R}^{N}:x_{i}\geq 0,\,\forall i=1,\ldots,N\}$. Then, we compute the MAP estimate using the proximal gradient descent algorithm in Eq.(31), and sample from the posterior using the MYULA algorithm in Eq.(32). The proximal operator of the positivity constraint, when applied to a vector $\mathbf{x}\in\mathbb{R}^{N}$, simply clips to zero all the negative entries of $\mathbf{x}$. This operation corresponds to a projection onto the positive orthant, which is a convex, closed set.

### v3

@@ 6 Summary, Conclusion and Next Steps
In this work, we discussed a unifying framework for Bayesian inference from plasma imaging tomographic data. We showed that many popular plasma emission tomographic techniques (such as isotropic/anisotropic approaches based on generalized Tikhonov or MFI regularization, Gaussian Process Tomography, and Maximum Likelihood tomography) can be naturally understood within a common Bayesian framework. Operating within such a framework, we leveraged well-studied optimization and inference concepts to propose and describe in detail a versatile pipeline that relies on simple, robust, and efficient algorithms. In particular, we proposed the unadjusted Langevin algorithm for efficient posterior sampling. We applied this pipeline to the problem of SXR imaging at TCV, demonstrating that the proposed approach enables principled, useful inference from noisy, sparse-view tomographic data. We validated the pipeline by conducting a phantom-based study over a large set of physically realistic SXR model phantoms. We showed that the pipeline leads to a statistically reliable estimation of the main features of the phantoms (in terms of shape, asymmetry, anisotropy, total SXR radiated power and peak location) and demonstrated the robustness of the approach with respect to moderate noise mismodeling and model mismatch. We showed that MAP computation can be complemented by MCMC methods for posterior sampling, which allow uncertainty quantification and complex statistical analysis but are more computationally expensive; such a complementary vision, with frameworks combining quick MAP-based analyses with detailed MCMC-based computations on data of particular interest, is becoming increasingly popular. Finally, we highlighted some of the inherent limitations of sparse-view tomography and the crucial role played by the prior in such an underdetermined setting: careful prior modeling is clearly essential. Moreover, we stressed the importance of adopting a probabilistic approach allowing to quantify the uncertainty associated to any estimate, which is made all the more vital by the sparse-view nature of plasma tomography. Open access is provided to the computational routines implemented in the context of this work, in an effort towards reproducibility and reusability: the link to the project repository can be found at the beginning of Section 5. Instructions are provided to generate the data used in this work, which are made available for further tomography studies. This software and data availability policy is chosen to facilitate building upon the proposed framework.
@@
For future work, we identify an applied and a methodological axis. From the application viewpoint, we will assess the scientific potential of the proposed pipeline by its application to tomographic reconstruction from the widest possible range of experimental TCV data. In particular, we will broaden the scope from SXR imaging to include bolometry and, potentially, visible-light imaging applications. This can be readily achieved within the generality of the proposed inference framework. Future studies could also investigate principled comparisons between techniques that have been typically viewed as distinct and unrelated, e.g., between traditional smoothness-promoting techniques and a fully Bayesian version of the ML method, as mentioned in Remark 4.0.1: this is now possible within the proposed unifying framework. From the methodological and modeling points of view, future studies could investigate priors that are different from the classical smoothness approaches traditionally considered for plasma imaging. Specifically, it would be very interesting to consider sparsity-promoting priors [32, 46] and recently popularized deep learning-based priors [62].
@@ A Phantom Analysis: supplementary results
$N_{1}$ $0.9\!\cdot\!10^{-2}$ $0.9\!\cdot\!10^{-2}$ $0.6\%$ $1.0\%$ 3.7cm3.7\,cm $4.4\%$ $1.1\%$
$N_{2}$ $1.3\!\cdot\!10^{-2}$ $3\!\cdot\!10^{-2} 1.1$$1\%$ $1.9\%$ 4.3cm4.3\,cm $5.9\%$ $1.6\%$
$N_{3}$ $1.2\!\cdot\!10^{-2}$ $1.3\!\cdot\!10^{-2}$ $0.9\%$ $1.4\%$ 4.6cm4.6\,cm $5.1\%$ $1.7\%$
@@ B Handling Non-Smooth Terms: Proximal Operators and Moreau-Yosida Approximations
$\log p(\mathbf{x}|\mathbf{y})=f(\mathbf{x})+g(\mathbf{x})\;,$ (26)
@@
In this work, we employ the reconstruction model in Eq.(24). While the exponential terms in Eq.(24) are smooth, the positivity constraint prior $\mathcal{P}_{\mathbf{x}\geq 0}=\exp(-\mathbbm{1}_{\mathbf{x}\geq 0})$ leads to a non-smooth, proximable term in the negative log-posterior. Therefore, we split $\log p(\mathbf{x}|\mathbf{y})$ into a smooth and non-smooth component as in Eq.(26), with $g=\mathbbm{1}_{\mathbf{x}\geq 0}$ defined as in Eq.(28) for $\mathcal{C}=\{\mathbf{x}\in\mathbb{R}^{N}:x_{i}\geq 0,\,\forall i=1,\ldots,N\}$. Then, we compute the MAP estimate using the proximal gradient descent algorithm in Eq.(31), and sample from the posterior using the MYULA algorithm in Eq.(32). The proximal operator of the positivity constraint, when applied to a vector $\mathbf{x}\in\mathbb{R}^{N}$, simply clips to zero all the negative entries of $\mathbf{x}$. This operation corresponds to a projection onto the positive orthant, which is a convex, closed set.